Develop have the (3. 15pts) Multidimensional arrays can be stored in row major order, as in C++. the access function for the row-major arrangement of three-dimensional arrays that following form: A[1b1. .ub1, 1b2. .ub2, 1b3. u and upper bounds for the respective dimensions of the array. Define the access function that a ub3], where lb and ubs represent the lower compiler needs to implement to compute the address of Al[i.j.K]. Address(ADJ.K]
Answer #1

Let the address of A[lb1,lb2,lb3] be base address.

the size of data type of array be S.

Address(A[i,j,k]) =

first of all,we need to reach ith 2-D.for that we need to cross or skip all the 2-D's that came before.(i.e i-lb1) 2-D's need to be crossed to reach ith 2-D.

Once we have reached ith 2-D,we need to reach jth row.for that we need to cross all rows that came before jth row.i.e( j-lb2 ) rows .

Once we need have reached desired row we need to reach desired column of jth row.to do that we need to skip or cross all columns before that.i.e.(k - lb3) columns.

Size of one 2-D = (ub2 - lb2 + 1)*(ub3 - lb3 + 1)*S

Size of one row = (ub3 - lb3 + 1)*S

Address(A[i,j,k]) =

Base address +( i-lb1)*(ub2-lb2+1)*(ub3 - lb3 + 1)*S + (j-lb2)*(ub3- lb3 + 1)*S + (k - lb3)*S